# Local analysis of iterative reconstruction from discrete generalized Radon transform data in the plane

### v1

@@ C.2. Extracting the leading local singularity of $f_{\epsilon}$
(D.12) $\sum_{j}\varphi(u-j)g(j)=g(u)+\textstyle\sum_{j}\varphi(u-j)R_{1}(j,u).$
Clearly
(D.13) $|g^{\prime\prime}(u)|=O\big{(}t^{-3/2}\big{)},\ \epsilon|t|<A,\ t\to\infty.$
Hence, by (D.11)
(D.14) $\begin{split}&R_{1}(j,u)=O(t^{-3/2}),\ |j-u|<A,\ \epsilon|t|<A,\ t\to\infty.\end{split}$
@@
(D.15) $\begin{split}\partial_{u}R_{1}(j,u)=&-g^{\prime\prime}(u)(j-u),\\ \partial_{u}\textstyle\sum_{j}\varphi(u-j)R_{1}(j,u)=&\textstyle\sum_{j}\varphi^{\prime}(u-j)R_{1}(j,u)-g^{\prime\prime}(u)\textstyle\sum_{j}\varphi(u-j)(j-u)\\ \hskip 56.9055pt=&\textstyle\sum_{j}\varphi^{\prime}(u-j)R_{1}(j,u),\end{split}$
@@
(D.16) $\begin{split}\partial_{u}^{2}\textstyle\sum_{j}\varphi(u-j)&R_{1}(j,u)\\ =&\textstyle\sum_{j}\varphi^{\prime\prime}(u-j)R_{1}(j,u)-g^{\prime\prime}(u)\textstyle\sum_{j}\varphi^{\prime}(u-j)(j-u)\\ \hskip 56.9055pt=&\textstyle\sum_{j}\varphi^{\prime\prime}(u-j)R_{1}(j,u)-g^{\prime\prime}(u).\end{split}$
@@
To prove (LABEL:H_props_2_a), by (6.10) and (6.11) we consider
@@
We moved one derivative inside the parentheses because $H$ is defined in terms of $h^{\prime}$ in (6.11). Replacing $\check{t}$ with $[\Phi(\theta+\epsilon\mu s)-\Phi(\theta)]/\epsilon+\check{t}+\check{q}$, multiplying both sides by $K(\theta+\epsilon\mu s)\varphi_{\alpha}(s)$, and integrating with respect to $s$ (cf. (6.10), (6.11)) finishes the proof. This also gives a formula for $\varkappa$. Our derivation shows that $\varkappa$ is smooth and bounded with all derivatives, but the formula for $\varkappa$ is not needed.
The claim (LABEL:H_props_2_b) immediately follows because $\varkappa$ is smooth and $\epsilon|\check{t}|<A$.
@@
For simplicity, in what follows we omit the dependence of $G(\check{t},\check{q};\theta,\epsilon)$ on $\check{q},\theta$, and $\epsilon$, and denote $Q^{\prime}:=Q_{\alpha}^{\prime}$ and $Q^{\prime\prime}:=Q_{\alpha\alpha}^{\prime\prime}$. By (6.15) we study the convergence of the sum
(D.19) $J_{\epsilon}:=\epsilon^{1/2}\sum_{\alpha_{j_{1}}\in\text{dom}\,P}G(Q(\alpha_{j_{1}})/\epsilon).$
Recall that $Q$ depends on $x$, which is confined to a sufficiently small open set $U^{\prime}$ (see the paragraph following (6.7)), and on the selected $P(\alpha)$.
If $|Q(\alpha)|\geq c$ on $\text{dom}\,P$, using (6.16) in (D.19) gives for each of the partial sums, over $Q(\alpha_{j_{1}})>0$ and $Q(\alpha_{j_{1}})<0$, $|J_{\epsilon}|\leq c\epsilon^{1/2}(1/\epsilon)\epsilon^{1/2}=O(1)$. Thus, in what follows we assume $Q(a_{0})=0$ for some $a_{0}\in\text{dom}\,P$ and replace $\text{dom}\,P$ in (D.19) with $I_{\alpha}^{\prime}$, which is a sufficiently small neighborhood of $a_{0}$. If $Q^{\prime}(\alpha)\not=0$ on $I_{\alpha}^{\prime}$, then $|Q^{\prime}(\alpha)|\asymp|\alpha-a_{0}|$ on $I_{\alpha}^{\prime}$ and
@@
Hence, in what follows, we assume that $Q^{\prime}(\tilde{\alpha})=0$ for some $\tilde{\alpha}\in I_{\alpha}^{\prime}$ and $\tilde{x}\in U^{\prime}$.
As is easily seen, $Q(\tilde{\alpha})=0$ and $Q^{\prime}(\tilde{\alpha})=0$ for some $\tilde{\alpha}\in\text{dom}\,P$ and $\tilde{x}\in U^{\prime}$ if and only if $\tilde{x}\in\mathcal{S}$ and $P(\alpha)$ is associated with $\tilde{x}$. The necessity is established in Lemma D.3 below, and the sufficiency follows from the Bolker condition, assumption 3.2(2). Recall that due to possible data redundancy, there can be several $P(\alpha)$ associated with $\tilde{x}$. In the rest of this subsection we select one of them. We need the following lemma, which is proven in appendix D.4.
@@
(D.21) $Q_{\alpha}^{\prime}(\tilde{x},\tilde{\alpha})=0,\ Q_{\alpha\alpha}^{\prime\prime}(\tilde{x},\tilde{\alpha})=\det M(\tilde{y})\frac{|x_{\alpha}^{\prime}(\tilde{\alpha})|^{2}}{|\Phi_{x}^{\prime}(\tilde{x},\tilde{\alpha})||H_{x}^{\prime}(\tilde{x})|}>0$
@@
(D.22) $|\Delta_{\Phi}(\tilde{x},\tilde{\alpha})|=|\det M(\tilde{y})|\frac{|x_{\alpha}^{\prime}(\tilde{\alpha})|}{|H_{x}^{\prime}(\tilde{x})|}.$
By continuity, $Q_{\alpha\alpha}^{\prime\prime}(x,\alpha)>0$ for all $(x,\alpha)\in U^{\prime}\times I_{\alpha}^{\prime}$.
D.3.1. Critical point in $I_{\alpha}^{\prime}$
Fix any $x\in U^{\prime}$. Suppose $Q^{\prime}(a_{1})=0$ for some $a_{1}\in I_{\alpha}^{\prime}$. By shifting $\alpha$, we can assume without loss of generality that $a_{1}=0$ and
(D.23) $Q(\theta)=-h+\int_{0}^{\theta}(\theta-s)Q^{\prime\prime}(s)\text{d}s$
for some $h$. The value of $h$ is determined from $Q(a_{0})=0$. Since $Q^{\prime\prime}(\alpha)\asymp 1$ on $I_{\alpha}^{\prime}$ (see the sentence following Lemma D.3) and $Q(a)=0$, we have $h\geq 0$.
@@ D.3.1. Critical point in $I_{\alpha}^{\prime}$
For any $c>0$, the number of $\alpha_{j_{1}}$ inside the set $\{\alpha\in I_{\alpha}^{\prime}:|Q(\alpha)|\leq c\epsilon\}$ is $O(\epsilon^{-1/2})$ and is uniformly bounded with respect to $a_{0}\in I_{\alpha}^{\prime}$. Since $G$ is bounded, the total contribution of such $\alpha_{j_{1}}$ to $J_{\epsilon}$ is bounded. Hence it remains to estimate the following two sums
(D.24) $\begin{split}J_{\epsilon}^{+}:=&\epsilon^{1/2}\sum_{\alpha_{j_{1}}\in I_{\alpha}^{\prime},Q(\alpha_{j_{1}})>\epsilon}\bigg{[}\epsilon^{1/2}+\frac{\epsilon}{Q(\alpha_{j_{1}})}\bigg{]},\\ J_{\epsilon}^{-}:=&\epsilon^{1/2}\sum_{\alpha_{j_{1}}\in I_{\alpha}^{\prime},Q(\alpha_{j_{1}})<-\epsilon}\bigg{(}-\frac{\epsilon}{Q(\alpha_{j_{1}})}\bigg{)}^{1/2}.\end{split}$
Begin with $J_{\epsilon}^{+}$. Clearly the sums over $\alpha_{j_{1}}>0$ and $\alpha_{j_{1}}<0$ lead to the same upper bound, so we consider only $\alpha_{j_{1}}>0$. Also, for estimation purposes, we can assume $Q^{\prime\prime}(s)\geq 1$, $s\in I_{\alpha}^{\prime}$ (rather than the more cumbersome $Q^{\prime\prime}(s)\geq c$). Let $\theta_{1}>0$ satisfy
(D.25) $Q(\theta_{1})=-h+\int_{0}^{\theta_{1}}(\theta_{1}-s)Q^{\prime\prime}(s)\text{d}s=\epsilon.$
For any $\theta\geq\theta_{1}$ one has
(D.26) $Q(\theta)\geq-h+\int_{0}^{\theta_{1}}(\theta_{1}-s)Q^{\prime\prime}(s)\text{d}s+\int_{\theta_{1}}^{\theta}(\theta-s)\text{d}s=\epsilon+(\theta-\theta_{1})^{2}/2.$
Since the shift $\theta\to\theta-\theta_{1}$ does not affect the upper bound we get
(D.27) $\begin{split}|J_{\epsilon}^{+}|\leq&c\epsilon^{1/2}\sum_{j_{1}=0}^{1/\epsilon}\bigg{[}\epsilon^{1/2}+\frac{\epsilon}{\epsilon+(\epsilon j_{1})^{2}}\bigg{]}\leq c+c\epsilon^{1/2}\int_{0}^{1}\frac{\text{d}x}{\epsilon+x^{2}}=O(1).\end{split}$
Next, consider $J_{\epsilon}^{-}$. If $h<\epsilon$, then $Q(\theta)\geq-\epsilon$ for all $\alpha\in I_{\alpha}^{\prime}$ and $J_{\epsilon}^{-}=0$. Hence we can assume $h>\epsilon$. In this case we initially assume for simplicity that $Q(0)=Q(a)=-\epsilon$ for some $a>0$. Therefore
(D.28) $\begin{split}-(Q(\theta)+\epsilon)=&-\int_{0}^{\theta}(\theta-s)Q^{\prime\prime}(s)\text{d}s+\theta\int_{0}^{a}(1-(s/a))Q^{\prime\prime}(s)\text{d}s\\ =&\int_{0}^{\theta}s(1-(\theta/a))Q^{\prime\prime}(s)\text{d}s+\theta\int_{\theta}^{a}(1-(s/a))Q^{\prime\prime}(s)\text{d}s.\end{split}$
Both integrands on the second line are positive, so the minimum on the right is achieved when $Q^{\prime\prime}(s)\equiv 1$. Hence, shifting the argument of $Q$ again we can assume $Q(\theta)=-h+\theta^{2}/2$. Similarly to (D.27),
(D.29) $\begin{split}|J_{\epsilon}^{-}|\leq&c\epsilon^{1/2}\sum_{h-(\epsilon j_{1})^{2}>\epsilon}\bigg{[}\frac{\epsilon}{h-(\epsilon j_{1})^{2}}\bigg{]}^{1/2}\leq c\int_{0}^{h^{1/2}}\frac{\text{d}x}{(h-x^{2})^{1/2}}<\infty.\end{split}$
Note that the value of the last integral is independent of $h>0$.
@@
It is clear that in the above arguments all the bounds can be made uniform with respect to $x\in U^{\prime}$.
@@
The first claim in (D.21) follows by differentiating (4.8) and (D.30) with respect to $\alpha$ and using that $x^{\prime}(\alpha)$ is tangent to $\mathcal{S}$, $x(\tilde{\alpha})=\tilde{x}$, and $\text{d}_{x}\Phi(\tilde{x},\tilde{\alpha})$ is conormal to $\mathcal{S}$ at $\tilde{x}$.
Differentiating (D.30) with respect to $\alpha$ twice and using (6.7) gives:
(D.31) $Q_{\alpha\alpha}^{\prime\prime}=-\big{[}(\Phi_{xx}^{\prime\prime}x_{\alpha}^{\prime},x_{\alpha}^{\prime})+2\Phi_{x\alpha}^{\prime\prime}x_{\alpha}^{\prime}+\Phi_{x}^{\prime}x_{\alpha\alpha}^{\prime\prime}\big{]}.$
Differentiate the second equation in (4.8) with respect to $\alpha$:
(D.32) $H_{xx}^{\prime\prime}x_{\alpha}^{\prime}=\nu_{1}^{\prime}\Phi_{x}^{\prime}+\nu_{1}\big{[}\Phi_{xx}^{\prime\prime}x_{\alpha}^{\prime}+\Phi_{x\alpha}^{\prime\prime}\big{]}.$
Compute the dot product of (D.32) with $x_{\alpha}^{\prime}$ and use that $\Phi_{x}^{\prime}x_{\alpha}^{\prime}\equiv H_{x}^{\prime}x_{\alpha}^{\prime}\equiv 0$ (which follows from (4.8) to get
(D.33) $(H_{xx}^{\prime\prime}x_{\alpha}^{\prime},x_{\alpha}^{\prime})=\nu_{1}\big{[}(\Phi_{xx}^{\prime\prime}x_{\alpha}^{\prime},x_{\alpha}^{\prime})+\Phi_{x\alpha}^{\prime\prime}x_{\alpha}^{\prime}\big{]}.$
Differentiate the identity $\Phi_{x}^{\prime}x_{\alpha}^{\prime}\equiv 0$ to get
(D.34) $(\Phi_{xx}^{\prime\prime}x_{\alpha}^{\prime},x_{\alpha}^{\prime})+\Phi_{x\alpha}^{\prime\prime}x_{\alpha}^{\prime}+\Phi_{x}^{\prime}x_{\alpha\alpha}^{\prime\prime}\equiv 0.$
Using (D.33) and (D.34) in (D.31) and applying Lemma 4.1 finally gives:
(D.35) $\begin{split}Q_{\alpha\alpha}^{\prime\prime}=&-\big{[}(1/\nu_{1})(H_{xx}^{\prime\prime}x_{\alpha}^{\prime},x_{\alpha}^{\prime})-(\Phi_{xx}^{\prime\prime}x_{\alpha}^{\prime},x_{\alpha}^{\prime})\big{]}\\ =&-\big{(}\big{[}H_{xx}^{\prime\prime}-\nu_{1}\Phi_{xx}^{\prime\prime}\big{]}e,e\big{)}\frac{|x_{\alpha}^{\prime}|^{2}}{\nu_{1}},\ e:=x_{\alpha}^{\prime}/|x_{\alpha}^{\prime}|.\end{split}$
Combining this with (B.1) and (B.2) proves the second claim in (D.21).
Next we prove (D.22). Similarly to (B.2),
(D.36) $|\Delta_{\Phi}(x,\alpha)|=|\Phi_{x}^{\prime}||\Phi_{x\alpha}^{\prime\prime}e|.$
Using (D.33) and then (B.2) gives
(D.37) $\begin{split}|\Phi_{x}^{\prime}||\Phi_{x\alpha}^{\prime\prime}e|&=\big{|}\big{(}[H_{xx}^{\prime\prime}-\nu_{1}\Phi_{xx}^{\prime\prime}]e,e\big{)}\big{|}\frac{|x_{\alpha}^{\prime}||\Phi_{x}^{\prime}|}{|\nu_{1}|}=\frac{|\text{det}M||x_{\alpha}^{\prime}|}{|\nu_{1}||\Phi_{x}^{\prime}|},\end{split}$

### v2

@@ C.2. Extracting the leading local singularity of $f_{\epsilon}$
(D.12) $g_{\epsilon}(u):=\textstyle\sum_{j}\varphi(u-j)g(j)=g(u)+\textstyle\sum_{j}\varphi(u-j)R_{1}(j,u).$
By assumption, $0<\epsilon t<A$ and $\epsilon|\check{q}|\leq A$, so $|\epsilon\check{t}+\epsilon\check{q}|<2A$ and
(D.13) $|g^{\prime\prime}(u)|=\left|\partial_{\check{t}}^{2}\left(f(\epsilon\check{t}+\epsilon\check{q})\check{t}^{1/2}\right)\right|=O\big{(}\check{t}^{-3/2}\big{)},\ \epsilon\check{t}<A,\ \check{t}\to+\infty.$
The values of $j,s,u$ used in (D.11) and (D.12) satisfy $|j-s|\leq|j-u|\leq\text{diam}(\text{supp}\,\varphi)$. Hence, by (D.11) and (D.13)
(D.14) $\begin{split}&R_{1}(j,u)=O(\check{t}^{-3/2}),\ j-u\in\text{supp}\,\varphi,\ \epsilon\check{t}<A,\ \check{t}\to+\infty.\end{split}$
@@
(D.15) $\begin{split}\partial_{u}R_{1}(j,u)=&-g^{\prime\prime}(u)(j-u),\\ g_{\epsilon}^{\prime}(u)-g^{\prime}(u)=&\partial_{u}\textstyle\sum_{j}\varphi(u-j)R_{1}(j,u)\\ =&\textstyle\sum_{j}\varphi^{\prime}(u-j)R_{1}(j,u)-g^{\prime\prime}(u)\textstyle\sum_{j}\varphi(u-j)(j-u)\\ \hskip 56.9055pt=&\textstyle\sum_{j}\varphi^{\prime}(u-j)R_{1}(j,u),\end{split}$
@@
(D.16) $\begin{split}g_{\epsilon}^{\prime\prime}(u)-g^{\prime\prime}(u)=&\textstyle\sum_{j}\varphi^{\prime\prime}(u-j)R_{1}(j,u)-g^{\prime\prime}(u)\textstyle\sum_{j}\varphi^{\prime}(u-j)(j-u)\\ \hskip 56.9055pt=&\textstyle\sum_{j}\varphi^{\prime\prime}(u-j)R_{1}(j,u)-g^{\prime\prime}(u).\end{split}$
@@
To prove (LABEL:H_props_2_a), in view of (6.13) and (6.14), we consider
@@
We moved one derivative inside the parentheses because $H$ is defined in terms of $h^{\prime}$ in (6.14). Replacing $\check{t}$ with $[\Phi(\theta+\epsilon\mu s)-\Phi(\theta)]/\epsilon+\check{t}$, multiplying both sides by $K(\theta+\epsilon\mu s)\varphi_{\alpha}(s)$, and integrating with respect to $s$ (cf. (6.13), (6.14)) finishes the proof. This also gives a formula for $\varkappa$:
(D.19) $\varkappa(t,q;\theta)=t\partial_{t}a_{0}(\theta,t+q)+(1/2)a_{0}(\theta,t+q).$
This shows that $\varkappa$ is smooth and bounded with all derivatives on any compact set. The claim (LABEL:H_props_2_b) immediately follows as well.
@@
D.3.1. Estimation away from the stationary points of $Q$.
It follows from (6.19) that the magnitude of $G$ can be controlled only by its first argument, $\check{t}$. Therefore, in this section, we omit the dependence of $G(\check{t},\check{q};*)$ on $\check{q}$ and $*=(x,\theta,\epsilon)$. In view of (6.18), we study the convergence of the sum
(D.20) $J_{\epsilon}:=\epsilon^{1/2}\textstyle\sum_{j_{1}\in J_{\alpha}}G(Q(x,\alpha_{j_{1}})/\epsilon),\ x\in U^{\prime}.$
Recall that $Q$ depends on the selected $P(\alpha)$.
Fix any $0<\delta\ll 1$ and $x\in U^{\prime}$. Since $x$ is fixed, we can drop it from notation. If $|Q(\alpha)|\geq\delta$ on $I_{\alpha}$, using (6.19) in (D.20) gives the same bound $|J_{\epsilon}|\leq c\epsilon^{1/2}(1/\epsilon)\epsilon^{1/2}=O(1)$ for each of the partial sums: over $Q(\alpha_{j_{1}})>0$ and over $Q(\alpha_{j_{1}})<0$. Thus, in what follows we may assume $|Q(\alpha)|\leq\delta$ for some $\alpha\in I_{\alpha}$. If $|Q_{\alpha}^{\prime}(\alpha)|\geq\delta$ on $I_{\alpha}$, then $|Q_{\alpha}^{\prime}(\alpha)|\asymp|\alpha-a_{0}|$ on $I_{\alpha}$ and
@@
D.3.2. Estimation near a stationary point of $Q$.
Based on the results of the preceding subsection, it remains to consider the case when $|Q(a_{1})|$, $|Q_{\alpha}^{\prime}(a_{2})|\leq\delta$ for some $a_{1},a_{2}\in I_{\alpha}$. Since $I_{\alpha}$ can be made as small as we like (but finitely small), it is instructive to consider the case when $Q(z,\alpha)=Q_{\alpha}^{\prime}(z,\alpha)=0$ for some $(z,\alpha)\in\mathcal{U}\times\mathcal{I}_{\alpha}$. This happens if and only if $z\in\mathcal{S}$ and $P(\alpha)$ is associated with $z$.
To prove sufficiency, suppose there is a point $(z,\alpha)\in\mathcal{U}\times\mathcal{I}_{\alpha}$ such that
(D.22) $\begin{split}&Q(z,\alpha)=\Phi(z,\alpha)-\Phi(x(\alpha),\alpha)=0,\\ &Q_{\alpha}^{\prime}(z,\alpha)=\Phi_{\alpha}^{\prime}(z,\alpha)-\Phi_{\alpha}^{\prime}(x(\alpha),\alpha)=0.\end{split}$
The Bolker condition, Assumption 3.2(2), gives $z=x(\alpha)$.
The necessity is established in the following Lemma D.3, which is proven in Appendix D.4.
@@
(D.23) $\begin{split}&Q(\tilde{x},\tilde{\alpha})=Q_{\alpha}^{\prime}(\tilde{x},\tilde{\alpha})=0,\\ &Q_{\alpha\alpha}^{\prime\prime}(\tilde{x},\tilde{\alpha})=\det M(\tilde{y})\frac{|x_{\alpha}^{\prime}(\tilde{\alpha})|^{2}}{|\text{d}_{x}\Phi(\tilde{x},\tilde{\alpha})||\text{d}_{x}H(\tilde{x})|}>0\end{split}$
@@
(D.24) $|\Delta_{\Phi}(\tilde{x},\tilde{\alpha})|=|\det M(\tilde{y})|\frac{|x_{\alpha}^{\prime}(\tilde{\alpha})|}{|\text{d}_{x}H(\tilde{x})|}.$
By the above argument, we can suppose that the point $x$ selected following (D.20) is located sufficiently close to the curve $x(\alpha),\alpha\in I_{\alpha}$. Otherwise, by shrinking $I_{\alpha}$ if necessary, we can ensure that either $|Q(\alpha)|\geq\delta$ or $|Q_{\alpha}^{\prime}(\alpha)|\geq\delta$ on $I_{\alpha}$. By Lemma D.3, $Q_{\alpha\alpha}^{\prime\prime}(x(\alpha),\alpha)\geq c>0$ for all $\alpha\in I_{\alpha}$. By continuity, $Q_{\alpha\alpha}^{\prime\prime}(\alpha)\geq c>0$ for all $\alpha\in I_{\alpha}$.
@@
Due to the lower bound on $Q_{\alpha\alpha}^{\prime\prime}$, the number of $\alpha_{j_{1}}$ in the set $\{\alpha\in I_{\alpha}:|Q(\alpha)|\leq c\epsilon\}$ is at most $A\epsilon^{-1/2}$, where $A$ is independent of $x$. Since $G$ is bounded, the total contribution of such $\alpha_{j_{1}}$ to $J_{\epsilon}$ is bounded. Hence, by (6.19), it remains to estimate the following two sums
(D.25) $\begin{split}J_{\epsilon}^{+}:=&\epsilon^{1/2}\sum_{\alpha_{j_{1}}\in I_{\alpha}^{+}}\bigg{[}\epsilon^{1/2}+\frac{\epsilon}{Q(\alpha_{j_{1}})}\bigg{]},\ I_{\alpha}^{+}:=\{\alpha\in I_{\alpha}:\,Q(\alpha)\geq\epsilon\},\\ J_{\epsilon}^{-}:=&\epsilon^{1/2}\sum_{\alpha_{j_{1}}\in I_{\alpha}^{-}}\bigg{(}-\frac{\epsilon}{Q(\alpha_{j_{1}})}\bigg{)}^{1/2},\ I_{\alpha}^{-}:=\{\alpha\in I_{\alpha}:\,Q(\alpha)\leq-\epsilon\}.\end{split}$
For estimation purposes, we can assume $Q^{\prime\prime}(\alpha)\geq 1$, $\alpha\in I_{\alpha}$ (rather than the more cumbersome $Q^{\prime\prime}(\alpha)\geq c$). When the argument $x$ of $Q$ is omitted and we write $Q(\alpha)$, then a prime denotes the derivative with respect to $\alpha$.
Since $Q^{\prime\prime}(\alpha)\geq 1$, $I_{\alpha}^{+}$ is the union of no more than two intervals, and $I_{\alpha}^{-}$ is a single interval. Also, the equation $Q^{\prime}(\alpha)=0$, $\alpha\in I_{\alpha}$, has at most one solution.
Begin with $J_{\epsilon}^{+}$. Let $[a,b]$ be one of at most three intervals that make up $I_{\alpha}^{+}$, subject to the condition $Q^{\prime}(\alpha)\not=0$ on $(a,b)$. Suppose $Q^{\prime}(\alpha)>0$ on $(a,b)$. The proof in the case $Q^{\prime}(\alpha)<0$ is analogous.
By construction, $Q(a)\geq\epsilon$. Consider the function $Q_{1}(\alpha)=\epsilon+(\alpha-a)^{2}/2$, $\alpha\in[a,b]$, which satisfies $Q_{1}(a)=\epsilon$, $Q_{1}^{\prime}(a)=0$, and $Q_{1}^{\prime\prime}(\alpha)=1$ on $[a,b]$. Clearly, $Q_{1}(\alpha)\leq Q(\alpha)$, $\alpha\in[a,b]$. Therefore, the value of $J_{\epsilon}^{+}$ is increased if we replace $Q$ with $Q_{1}$. There are $O(1/\epsilon)$ terms in the first sum in (D.25), so the term $O(\epsilon^{1/2})$ in the brackets in $J_{\epsilon}^{+}$ contributes an $O(1)$ term. Then
(D.26) $\begin{split}J_{\epsilon}^{+}\leq&c+c\epsilon^{1/2}\sum_{j}\frac{\epsilon}{\epsilon+(\epsilon j)^{2}/2}\leq c.\end{split}$
Next, consider $J_{\epsilon}^{-}$. Let $[a,b]$ be one of at most two intervals that make up $I_{\alpha}^{-}$, subject to the condition $Q^{\prime}(\alpha)\not=0$ on $(a,b)$. Again, we can assume $Q^{\prime}(\alpha)>0$ on $(a,b)$. Consider the function $Q_{1}(\alpha)=Q(a)+(\alpha-a)^{2}/2$, $\alpha\geq a$, which satisfies $Q_{1}(a)=Q(a)$, $Q_{1}^{\prime}(a)=0$, and $Q_{1}^{\prime\prime}(\alpha)=1$ for $\alpha\geq a$ (see the green curve in Fig. 5). By construction, $Q_{a}^{\prime}(\alpha)\leq Q^{\prime}(\alpha)$, $\alpha\in[a,b]$. Pick any $t\in[a,b]$, and find $t_{1}$ by solving $Q_{1}(t_{1})=Q(t)$, i.e. $t_{1}=a+[2(Q(t)-Q(a))]^{1/2}$ (see Fig. 5). We claim that $Q_{1}^{\prime}(t_{1})\leq Q^{\prime}(t)$.
Figure 5. Construction used to estimate $J_{\epsilon}^{-}$.
To prove the claim, we need to show that $t_{1}-a=[2(Q(t)-Q(a))]^{1/2}<Q^{\prime}(t)$. This is equivalent to
(D.27) $2\bigg{[}Q^{\prime}(a)t+\int_{a}^{t}(t-s)Q^{\prime\prime}(s)\text{d}s\bigg{]}\leq\bigg{[}Q^{\prime}(a)+\int_{a}^{t}Q^{\prime\prime}(s)\text{d}s\bigg{]}^{2},\ a\leq t\leq b.$
Squaring the right-hand side and using that $Q^{\prime}(a)\geq 0$ and $Q^{\prime\prime}(s)\geq 1$, we see that the desired result follows if we prove
(D.28) $2\int_{a}^{t}(t-s)Q^{\prime\prime}(s)\text{d}s\leq\bigg{[}\int_{a}^{t}Q^{\prime\prime}(s)\text{d}s\bigg{]}^{2}.$
Writing $Q^{\prime\prime}(s)=1+\phi(s)$, where $\phi(s)\geq 0$, and simplifying proves the claim.
Let $b_{1}$ be obtained by solving $Q_{1}(b_{1})=Q(b)$. We have $Q_{1}^{\prime}(\alpha)\leq Q^{\prime}(\alpha)$, so $b_{1}\geq b$. The preceding argument implies that if the graph of $Q_{1}$ is shifted by $b_{1}-b$ to the left (to obtain the blue curve in Fig. 5 as indicated by the black arrow), then it will lie above the graph of $Q$ (shown in red). This means that $Q_{1}(\alpha+(b_{1}-b))\geq Q(\alpha)$, $\alpha\in[a,b]$. Hence, if we replace $Q(\alpha)$ by $Q_{1}(\alpha+(b_{1}-b))$ and $[a,b]$ by $[a-(b_{1}-b),b]$, then the value of $J_{\epsilon}^{-}$ is increased. The shift, of course, is irrelevant for estimation purposes, so we instead consider $Q_{1}(\alpha)$ and $[a,b_{1}]$. This leads to the following bound
(D.29) $\begin{split}J_{\epsilon}^{-}&\leq c\epsilon^{1/2}\sum_{h+(\epsilon j)^{2}/2<-\epsilon}\bigg{[}\frac{\epsilon}{-h-(\epsilon j)^{2}/2}\bigg{]}^{1/2}\leq c\int_{0}^{|2h|^{1/2}}\frac{\text{d}x}{(|2h|-x^{2})^{1/2}}<\infty.\end{split}$
Note that the value of the last integral is independent of $h$.
@@
It is clear that in the above arguments all the bounds are uniform with respect to $x\in U^{\prime}$.
@@
By construction, $x^{\prime}(\alpha)$ is tangent to $\mathcal{S}$ and $\text{d}_{x}\Phi(x(\alpha),\alpha)$ is conormal to $\mathcal{S}$ at $x(\alpha)$ (see (4.8)). Differentiation of (D.30) with respect to $\alpha$ gives
(D.31) $\begin{split}Q^{\prime}(\alpha)=&\Phi_{\alpha}^{\prime}(\tilde{x},\alpha)-\big{[}\text{d}_{x}\Phi(x(\alpha),\alpha)\cdot x_{\alpha}^{\prime}+\Phi_{\alpha}^{\prime}(x(\alpha),\alpha)\big{]}\\ =&\Phi_{\alpha}^{\prime}(\tilde{x},\alpha)-\Phi_{\alpha}^{\prime}(x(\alpha),\alpha).\end{split}$
Using that $x(\tilde{\alpha})=\tilde{x}$ proves the first line in (D.23).
Next, differentiate (D.31) with respect to $\alpha$ and set $\alpha=\tilde{\alpha}$ to obtain
(D.32) $Q_{\alpha\alpha}^{\prime\prime}=-\text{d}_{x}\Phi_{\alpha}^{\prime}\cdot x_{\alpha}^{\prime}.$
Differentiate the second equation in (4.8) with respect to $\alpha$ and set $\alpha=\tilde{\alpha}$:
(D.33) $\text{d}_{x}^{2}H\,x_{\alpha}^{\prime}=\nu_{1}^{\prime}\text{d}_{x}\Phi+\nu_{1}\big{[}\text{d}_{x}^{2}\Phi\,x_{\alpha}^{\prime}+\text{d}_{x}\Phi_{\alpha}^{\prime}\big{]}.$
Apply (D.33) to $x_{\alpha}^{\prime}$ and use that $\text{d}_{x}\Phi\cdot x_{\alpha}^{\prime}\equiv 0$ a second time to get
(D.34) $\text{d}_{x}^{2}H(x_{\alpha}^{\prime},x_{\alpha}^{\prime})=\nu_{1}\big{[}\text{d}_{x}^{2}\Phi(x_{\alpha}^{\prime},x_{\alpha}^{\prime})+\text{d}_{x}\Phi_{\alpha}^{\prime}\cdot x_{\alpha}^{\prime}\big{]}.$
Using (D.34) in (D.32) and applying Lemma 4.1 finally gives:
(D.35) $\begin{split}Q_{\alpha\alpha}^{\prime\prime}=&-\big{[}(1/\nu_{1})\text{d}_{x}^{2}H(x_{\alpha}^{\prime},x_{\alpha}^{\prime})-\text{d}_{x}^{2}\Phi(x_{\alpha}^{\prime},x_{\alpha}^{\prime})\big{]}\\ =&-\text{d}_{x}^{2}[H-\nu_{1}\Phi](e,e)\frac{|x_{\alpha}^{\prime}|^{2}}{\nu_{1}},\ e:=x_{\alpha}^{\prime}/|x_{\alpha}^{\prime}|.\end{split}$
Combining this with (B.1) and (B.4) proves the second line in (D.23).
Next we prove (D.24). By (3.4), (3.5), $\Delta_{\Phi}(x,\alpha)=\det\begin{pmatrix}\text{d}_{x}\Phi\\ \text{d}_{x}\Phi_{\alpha}^{\prime}\end{pmatrix}$. Considering the matrix product $\Delta_{\Phi}(x,\alpha)V$, where $V=\begin{pmatrix}e_{0}&e\end{pmatrix}$ (see (B.2)), gives
(D.36) $|\Delta_{\Phi}(x,\alpha)|=|\text{d}_{x}\Phi||\text{d}_{x}\Phi_{\alpha}^{\prime}\cdot e|.$
Using (D.32), (D.35), and then (B.4) gives
(D.37) $\begin{split}|\text{d}_{x}\Phi||\text{d}_{x}\Phi_{\alpha}^{\prime}\cdot e|&=\big{|}\text{d}_{x}^{2}[H-\nu_{1}\Phi](e,e)\big{|}\frac{|\text{d}_{x}\Phi||x_{\alpha}^{\prime}|}{|\nu_{1}|}=\frac{|\text{det}M||x_{\alpha}^{\prime}|}{|\nu_{1}||\text{d}_{x}\Phi|},\end{split}$